9-59 Two equipment investments are estimated as follows:

Year A B
0 -15,000 -18,000
1 5,000 6,500
2 5,000 6,500
3 5,000 6,500
4 5,000 6,500
5 5,000 6,500

Which investment has the better discounted payback period if interest = 14%
Answer #1

Equipment A:

Col 1 Col 2 Col 3 Col 4 Col 5
Year Cash flows ($) Present Value Factor (P/F, 14%, n) Discounted Cash Flow Col 2 * Col 3 Cumulative Discounted cash flows($)
0 -15000 1.0000 -15000.00 -15000.00
1 5000 0.8772 4386.00 -10614.00
2 5000 0.7695 3847.50 -6766.50
3 5000 0.6750 3375.00 -3391.50
4 5000 0.5921 2960.50 -431.00
5 5000 0.5194 2597.00 2166.00

Discounted payback period = 4 + |-431| / 2597 = 4 + (431 / 2597) = 4 + 0.17 = 4.17 years

Equipment B:

Col 1 Col 2 Col 3 Col 4 Col 5
Year Cash flows ($) Present Value Factor (P/F, 14%, n) Discounted Cash Flow Col 2 * Col 3 Cumulative Discounted cash flows($)
0 -18000 1.0000 -18000.00 -18000.00
1 6500 0.8772 5701.80 -12298.20
2 6500 0.7695 5001.75 -7296.45
3 6500 0.6750 4387.50 -2908.95
4 6500 0.5921 3848.65 939.70
5 6500 0.5194 3376.10 4315.80

Discounted payback period = 4 + |-939.70| / 3376.10 = 4 + (939.70 / 3376.10) = 4 + 0.28 = 4.28 years

Thus, investment A has the better discounted payback period.